OpenStudy (anonymous):

If y varies directly as x, and y=8 when x=6 , find x when y=16?

OpenStudy (anonymous):

y = kx find value of k by plugging in x=6 y=8 then you can find x when y = 16

OpenStudy (hoblos):

y=8 when x=6 then y=8/6 x = 4/3 x for y=16 => x= 3/4y= 12

OpenStudy (anonymous):

yup